The probability that a student drinks coffee is 0.50; the probability that a student ride a bus to school is 0.30; the probability that a student does both is 0.20. Therefore, the frequency of drinking coffee and riding a bus to school are independent of one another. True Or False
Answer #1

FALSE.

The events A and B are independent if P(A ∩ B) = P(A) P(B).

Let A and B be two events i.e the student drinks coffee and a student rides a bus to school.

P(A B) = 0.2 i,e the Students does both

But, P(A).P(B) = 0.5* 0.3 = 0.15

Therefore, 0.15 0.2

A and B are DEPENDENT.